Each edge of a cube is colored either red or black. If every face of the cube has at least one black edge, what is the smallest possible number of black edges?
Answer by richwmiller(17219)

There are 6 faces. Each face has 4 edges but not unique edges since edges are shared with one other face. 6*4=24 24/2=12 unique edges. So at least one of four edges is black. Each black edge is shared. So we need at least 3 black edges to have all faces with a black edge. 12/4=3